OpenStudy (anonymous):

find the length of the parametric curve. x=2t y=t ^{3}/3+1/t 1

OpenStudy (anonymous):

\[\int\limits_{1}^{5}\sqrt{(dx/dt)^{2} + (dy/dt)^{2}}dt\]

OpenStudy (anonymous):

can you show me the step? :(

OpenStudy (anonymous):

\[\int\limits_{1}^{5}\sqrt{4 + (t ^{2} - 1/t ^{2})^{2}}dt\]

OpenStudy (anonymous):

the part under the root becomes (t^2 + t^(-2))^2

OpenStudy (anonymous):

so its t^2dt + t^(-2)dt

OpenStudy (anonymous):

so thats the final answer??

OpenStudy (watchmath):

jojo, you need to integrate him1618 last answer from 1 to 5.

OpenStudy (anonymous):

okay thanks for the hint! :)